Two small conducting balls of identical mass m and charge q hang from non-conducting rope of length L Assume that Ɵ is very small tanƟ = sinƟ. Find an expression for the equilibrium separation x of the balls.


1
Expert's answer
2021-04-26T16:59:56-0400

Answer

Tension in rope

By vertical equilibrium

"T=\frac{mg}{cos\theta}"

By equilibrium

"\frac{kq^2}{L^2sin^2\theta}=Tsin{\theta}"


And

"\frac{kq^2}{L^2sin^2\theta}=\frac{mg}{cos\theta}sin{\theta}"


"\frac{kq^2}{4(L^2sin^2\theta)}=mgtan{\theta}"